OpenStudy (anonymous):

A data set is shown. 48, 43, 42, 49, 50, 41 What is the mean absolute deviation of this data set? 45.5 9 3.5 3

jimthompson5910 (jim_thompson5910):

first off, what is the mean of this data set?

OpenStudy (anonymous):

15.25

jimthompson5910 (jim_thompson5910):

no you made an error somewhere

OpenStudy (anonymous):

45.5

jimthompson5910 (jim_thompson5910):

yep that's the mean but not the mean absolute deviation

jimthompson5910 (jim_thompson5910):

what you do next is subtract the mean from each data value then apply absolute value bars like this... |48 - 45.5| = 2.5 |43 - 45.5| = 2.5 |42 - 45.5| = 3.5 |49 - 45.5| = 3.5 |50 - 45.5| = 4.5 |41 - 45.5| = 4.5

OpenStudy (anonymous):

then add them together?

jimthompson5910 (jim_thompson5910):

correct, then you average those results (add them up, then divide by 6)

OpenStudy (anonymous):

ahh thank you!

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

so 3.5?

jimthompson5910 (jim_thompson5910):

yep 2.5 + 2.5 + 3.5 + 3.5 + 4.5 + 4.5 = 21 21/6 = 3.5